# Sampling-based Distributed Training with Message Passing Neural Network

### v3

@@ 1 Introduction and Related Work
The application of machine learning algorithms to build surrogates for partial differential equations (PDEs) has seen a significant push in recent years Li et al. (2020b); Raissi et al. (2019); Lu et al. (2021); Pfaff et al. (2020). While the majority of these ML models for PDEs surrogate modeling are CNN-based (Zhu et al., 2019; Ranade et al., 2022; Ren et al., 2022), the introduction of graph neural networks (Kipf & Welling, 2017) has made graph-based modeling increasingly common. Traditional numerical methods for real-world applications predominantly rely on mesh-based unstructured representations of the computational domains Mavriplis (1997) owing to the ease of (a) discretization of complex domains and (b) extension to adaptive mesh refinement strategies to capture multi-scale physical phenomenon like turbulence, boundary layers in fluid dynamics, etc. These mesh-based unstructured formulations can be seamlessly formulated on graphs. Furthermore, a graph-based representation eliminates the need for interpolating data onto a structured Euclidean grid – a step inherent in CNNs, further reducing cost and error. In recent years, these graph-based models have also been applied to crucial domains like weather modeling (Lam et al., 2022), biomedical science (Fout et al., 2017), computational chemistry (Gilmer et al., 2017; Duvenaud et al., 2015), etc.
From the perspective of physical system dynamics modeling, we can divide the GNN-based methods into two broad categories: (a) node-based approaches and (b) edge-based approaches. The node-based approaches, including graph convolution networks (GCNs)  (Kipf & Welling, 2017), GraphSAGE  (Hamilton et al., 2017), and graph attention networks (GATs)  (Veličković et al., 2018), are easier to scale with an increasing number of elements in an unstructured grid setting. However, they lack the mechanism of message-passing among nodes which is crucial to the modeling of spatial relationships between the nodes in complex PDEs. Hence, the edge-based graph methods dominate the paradigm of graph-based modeling for physical systems, such as Graph Network-based Simulators (Sanchez-Gonzalez et al., 2020), and its evolution MeshGraphNets (MGN) (Pfaff et al., 2020).  Pfaff et al. (2020) indicate that these edge-based graph methods outperform node-based GCNs in accuracy and stability, and can be applied to a wide variety of simulations. Another approach to edge-based methods for PDE modeling is message-passing neural networks (Gilmer et al., 2017; Li et al., 2020a) where the input and the output solution space are invariant to mesh grids and independent of discretization (Li et al., 2020a). Message-passing neural networks (MPNN) (Simonovsky & Komodakis, 2017) condition weights around the vertices based on edge attributes and have properties to learn the PDE characteristics from a sparsely sampled field.
@@ 2 Model Description
The message-passing graph model used in this work uses a convolutional graph neural network with edge conditioning (Simonovsky & Komodakis, 2017) together with random sampling (Li et al., 2020a; Bonnet et al., 2022b).
@@ 2.1 Graph construction
At the core of this model is the concept of using edge-conditioned convolution (Simonovsky & Komodakis, 2017) to calculate the node attribute $v_{i}^{l}$ in one message-passing step by summing the product of weights $\mathcal{K}_{\phi}$, based on the edge attributes $e_{ij}$ and neighboring node attributes $v_{j}$ in $\mathcal{G}_{i}$, giving
@@ 3 Methodology
In the multi-GPU setup, domains are equally partitioned based on their coordinates, and a distributed communication package from PyTorch facilitates inter-node communication (Paszke et al., 2019). The neural network architecture employed in this study consists of a 3-layer encoder and a 3-layer decoder, augmented by a convolution kernel. The network, with approximately $700$k parameters, is utilized consistently across all experiments involving DS-MPNN and the baseline Graph Convolution Networks (GCN) This uniformity ensures parameter consistency across different models, and all models are implemented using PyTorch geometric package (Fey & Lenssen, 2019). GCN here uses 6 hidden layers with a size of 378. Post-encoding, the latent space representation for all experiments apart from parametric studies maintains a consistent dimensionality of 32 attributes. Optimization is conducted using the Adam optimizer, complemented by the OneCycleLR scheduler (Smith & Topin, 2019), to enhance learning efficiency and effectiveness.
@@ 5 Conclusions
We address the issue of scaling edge-based graph methods to larger numbers of nodes by introducing distributed training for message-passing neural networks (MPNNs). When combined with node sampling techniques, this distributed approach allows us to scale a larger number of nodes with no or minimal loss in accuracy as compared to the single-GPU implementation while achieving a considerable decrease in training and inference times. We also show comparisons to the graph convolution networks (GCNs) and establish that edge-based methods like MPNNs outperform node-based methods like GCNs. This work opens up new avenues for the use of edge-based graph neural networks in problems of practical interest where the number of nodes can be impractically large, $\sim O(10^{5}-10^{6})$, for single-GPU memory limits. The limitations of DS-MPNN lie in the complexity of managing unstructured grids and adjacent elements. However, distributed graph training is advancing rapidly, with new libraries such as PyTorch Geometric version 2.5.0 (Fey & Lenssen, 2019) addressing these challenges. Future work should also involve the use of more sophisticated partitioning methods employed in engineering simulations, such as METIS.

### v4

@@ 1 Introduction and Related Work
The application of machine learning algorithms to build surrogates for partial differential equations (PDEs) has seen a significant push in recent years Li et al. (2020b); Raissi et al. (2019); Lu et al. (2021); Pfaff et al. (2020). While the majority of these ML models for PDEs surrogate modeling are CNN-based (Zhu et al., 2019; Ranade et al., 2022; Ren et al., 2022), the introduction of graph neural networks (Kipf and Welling, 2017) has made graph-based modeling increasingly common. Traditional numerical methods for real-world applications predominantly rely on mesh-based unstructured representations of the computational domains Mavriplis (1997) owing to the ease of (a) discretization of complex domains and (b) extension to adaptive mesh refinement strategies to capture multi-scale physical phenomenon like turbulence, boundary layers in fluid dynamics, etc. These mesh-based unstructured formulations can be seamlessly formulated on graphs. Furthermore, a graph-based representation eliminates the need for interpolating data onto a structured Euclidean grid – a step inherent in CNNs, further reducing cost and error. In recent years, these graph-based models have also been applied to crucial domains like weather modeling (Lam et al., 2022), biomedical science (Fout et al., 2017), computational chemistry (Gilmer et al., 2017; Duvenaud et al., 2015), etc.
From the perspective of physical system dynamics modeling, we can divide the GNN-based methods into two broad categories: (a) node-based approaches and (b) edge-based approaches. The node-based approaches, including graph convolution networks (GCNs)  (Kipf and Welling, 2017), GraphSAGE  (Hamilton et al., 2017), and graph attention networks (GATs)  (Veličković et al., 2018), are easier to scale with an increasing number of elements in an unstructured grid setting. However, they lack the mechanism of message-passing among nodes which is crucial to the modeling of spatial relationships between the nodes in complex PDEs. Hence, the edge-based graph methods dominate the paradigm of graph-based modeling for physical systems, such as Graph Network-based Simulators (Sanchez-Gonzalez et al., 2020), and its evolution MeshGraphNets (MGN) (Pfaff et al., 2020).  Pfaff et al. (2020) indicate that these edge-based graph methods outperform node-based GCNs in accuracy and stability, and can be applied to a wide variety of simulations. Another approach to edge-based methods for PDE modeling is message-passing neural networks (Gilmer et al., 2017; Li et al., 2020a) where the input and the output solution space are invariant to mesh grids and independent of discretization (Li et al., 2020a). Message-passing neural networks (MPNN) (Simonovsky and Komodakis, 2017) condition weights around the vertices based on edge attributes and have properties to learn the PDE characteristics from a sparsely sampled field.
@@ 2 Model Description
The message-passing graph model used in this work uses a convolutional graph neural network with edge conditioning (Simonovsky and Komodakis, 2017) together with random sampling (Li et al., 2020a; Bonnet et al., 2022b).
@@ 2.1 Graph construction
At the core of this model is the concept of using edge-conditioned convolution (Simonovsky and Komodakis, 2017) to calculate the node attribute $v_{i}^{l}$ in one message-passing step by summing the product of weights $\mathcal{K}_{\phi}$, based on the edge attributes $e_{ij}$ and neighboring node attributes $v_{j}$ in $\mathcal{G}_{i}$, giving
@@ 3 Methodology
In the multi-GPU setup, domains are equally partitioned based on their coordinates, and a distributed communication package from PyTorch facilitates inter-node communication (Paszke et al., 2019). The neural network architecture employed in this study consists of a 3-layer encoder and a 3-layer decoder, augmented by a convolution kernel. The network, with approximately $700$k parameters, is utilized consistently across all experiments involving DS-MPNN and the baseline Graph Convolution Networks (GCN) This uniformity ensures parameter consistency across different models, and all models are implemented using PyTorch geometric package (Fey and Lenssen, 2019). GCN here uses 6 hidden layers with a size of 378. Post-encoding, the latent space representation for all experiments apart from parametric studies maintains a consistent dimensionality of 32 attributes. Optimization is conducted using the Adam optimizer, complemented by the OneCycleLR scheduler (Smith and Topin, 2019), to enhance learning efficiency and effectiveness.
@@ 5 Conclusions
We address the issue of scaling edge-based graph methods to larger numbers of nodes by introducing distributed training for message-passing neural networks (MPNNs). When combined with node sampling techniques, this distributed approach allows us to scale a larger number of nodes with no or minimal loss in accuracy as compared to the single-GPU implementation while achieving a considerable decrease in training and inference times. We also show comparisons to the graph convolution networks (GCNs) and establish that edge-based methods like MPNNs outperform node-based methods like GCNs. This work opens up new avenues for the use of edge-based graph neural networks in problems of practical interest where the number of nodes can be impractically large, $\sim O(10^{5}-10^{6})$, for single-GPU memory limits. The limitations of DS-MPNN lie in the complexity of managing unstructured grids and adjacent elements. However, distributed graph training is advancing rapidly, with new libraries such as PyTorch Geometric version 2.5.0 (Fey and Lenssen, 2019) addressing these challenges. Future work should also involve the use of more sophisticated partitioning methods employed in engineering simulations, such as METIS.